# Preventing Model Collapse Under Overparametrization: Optimal Mixing Ratios for Interpolation Learning and Ridge Regression

### v1

@@ 4 Simulations
Figure 2: In panels $(a)$ and $(b)$, we plot the optimal mixing weight $w^{\star}$ as a function of $\lambda$ for different values of $\gamma$ and two classes of covariance matrices:. Panel (a) considers isotropic covariance ${\bm{\Sigma}}=I$ with high noise $\sigma^{2}=64$, demonstrating $w^{\star}$ can be close to $0.5$ for low SNR. Panel $(b)$ plots $w^{\star}$ for the spiked covariance matrix showing $w^{\star}=1$ for large $\lambda$. Panel $(c)$ plots the generalization error as a function of $\lambda$ for $\bm{\Sigma}$} equicorrelated matrix. Here, empirical risks align with theoretical predictions given by Proposition 3.1, even though $\bm{\Sigma}$} violates Assumption 2.1, illustrating the robustness of our results.
@@ 5 Discussion
We provide a rigorous analysis of model collapse under overparametrization for linear models. As an overarching theme, we demonstrate how mixing real-data with synthetic outputs mitigates model collapse, and identify optimal mixing ratios that minimize prediction error in this context. As a promising next direction, understanding how model collapse affects interpolators in other $\ell_{p}$ geometries would be crucial. Such interpolators arise as implicit regularized limits of popular algorithms Gunasekar et al. (2018b) and typically require techniques beyond random matrix theory—a critical tool employed in our analysis. Additionally, approaches that extend linear arguments to non-linear high-dimensional problems (c.f., Sur & Candès (2019); Hu & Lu (2022)) should enable our qualitative conclusions to generalize to structured non-linear models. This includes generalized linear models, single-index models, and even non-parametric models through parametric-to-non-parametric equivalence techniques introduced in Lahiry & Sur (2023). Finally, extending our results to more complex architectures remains an important challenge. One promising approach involves studying multi-index models and their sequential variants, which capture classes of neural networks and transformer architectures Troiani et al. (2024; 2025); Cui (2025). Investigating model collapse and mitigation strategies in these contexts presents an exciting avenue for future research.
@@ A.1 Notations
$M:=\mathbf{X}^{\top}\mathbf{X},\quad\quad A_{\lambda}=(M+n\lambda I)^{-1}.$
Note that $A_{\lambda}$ and $M$ are simultaneously diagonalizable and thus commute with each other. If $\lambda=0$, we use the notation $A_{0}=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$. The mixing proportion of the real-data is denoted by $w$ and we define $\widetilde{w}:=1-w$. For two sequence of random variables $u_{n},v_{n}$, we use the notation $u_{n}\sim v_{n}$ if $u_{n}/v_{n}\xrightarrow{P}1$ as $n\rightarrow\infty$.
@@ A.2 Variance
$\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2(k-1)}M^{2k-1}A_{\lambda}^{2k}+\widetilde{w}^{2t}M^{2t+1}A_{\lambda}^{2t+2}\right]$ (A.1)
@@
We will prove equation (A.1) by induction. For the base case $(t=0)$, we have $\widehat{\boldsymbol{\beta}}_{0,\lambda}=A_{\lambda}\mathbf{X}^{\top}\mathbf{y}$. Hence,
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{0,\lambda}|\mathbf{X}]$ $=A_{\lambda}\mathbf{X}^{\top}{\sf Cov}[\mathbf{y}|\mathbf{X}]\mathbf{X}A_{\lambda}$
$=A_{\lambda}\mathbf{X}^{\top}(\sigma^{2}I)\mathbf{X}A_{\lambda}=\sigma^{2}A_{\lambda}^{2}M$
@@
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t+1,\lambda}|\mathbf{X}]$ $=A_{\lambda}\mathbf{X}^{\top}{\sf Cov}[w\mathbf{y}_{t}+\widetilde{w}\mathbf{y}_{t}|\mathbf{X}]\mathbf{X}A_{\lambda}$
$=A_{\lambda}\mathbf{X}^{\top}(w^{2}\sigma^{2}I+\widetilde{w}^{2}\mathbf{X}{\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]\mathbf{X}^{\top}+\sigma^{2}\widetilde{w}^{2}I)\mathbf{X}A_{\lambda}$
$=\widetilde{w}^{2}A_{\lambda}M{\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]MA_{\lambda}+\sigma^{2}(w^{2}+\widetilde{w}^{2})MA_{\lambda}^{2}$
$=\widetilde{w}^{2}A_{\lambda}M\left(\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2(k-1)}M^{2k-1}A_{\lambda}^{2k}+\widetilde{w}^{2t}M^{2t+1}A_{\lambda}^{2t+2}\right]\right)MA_{\lambda}$
$\quad\quad+\sigma^{2}(w^{2}+\widetilde{w}^{2})MA_{\lambda}^{2}$
$=\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2k}M^{2k+1}A_{\lambda}^{2k+2}+\widetilde{w}^{2t+2}M^{2t+3}A_{\lambda}^{2t+4}\right]$
$\quad\quad+\sigma^{2}(w^{2}+\widetilde{w}^{2})MA_{\lambda}^{2}$
$=\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t+1}\widetilde{w}^{2(k-1)}M^{2k-1}A_{\lambda}^{2k}+\widetilde{w}^{2t+2}M^{2t+3}A_{\lambda}^{2t+4}\right]$
@@
$V(\boldsymbol{\beta}_{t,\lambda};\boldsymbol{\beta})$ $=\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2(k-1)}\operatorname{Tr}[B_{\lambda}^{k}]+\widetilde{w}^{2t}\operatorname{Tr}[B_{\lambda}^{t+1}]\right]$ (A.2)
where $B_{\lambda}^{k}=M^{2k-1}A_{\lambda}^{2k}{\bm{\Sigma}}$
Next, note that $\sigma_{\max}(MA_{\lambda})<1$ and therefore ${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]$ converges in matrix norm to the following limit:
$\lim_{t\to\infty}{\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]$ $=\sigma^{2}(w^{2}+(1-w)^{2})\sum_{k=0}^{\infty}(1-w)^{2k}(\mathbf{X}^{\top}\mathbf{X})^{2k+1}A_{\lambda}^{2k+2}$
$=\sigma^{2}(w^{2}+(1-w)^{2})\mathbf{X}^{\top}\mathbf{X}A_{\lambda}^{2}(I-(1-w)^{2}(\mathbf{X}^{\top}\mathbf{X})^{2}A_{\lambda}^{2})^{-1}$
@@
$=\frac{\sigma^{2}(w^{2}+(1-w)^{2})}{2(1-w)}\left[(w\mathbf{X}^{\top}\mathbf{X}+n\lambda I)^{-1}-((2-w)\mathbf{X}^{\top}\mathbf{X}+n\lambda I)^{-1}\right]$
$\implies\lim_{t\to\infty}V(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\frac{\sigma^{2}(w^{2}+(1-w)^{2})}{2(1-w)}\left(\frac{1}{w}\operatorname{Tr}[A_{\lambda/w}{\bm{\Sigma}}]-\frac{1}{2-w}\operatorname{Tr}[A_{\lambda/(2-w)}{\bm{\Sigma}}]\right)$
@@
$\lim_{n\to\infty}\operatorname{Tr}[A_{-z}{\bm{\Sigma}}]=\lim_{n\to\infty}\frac{1}{n}\operatorname{Tr}\left[(n^{-1}\mathbf{X}^{\top}\mathbf{X}-zI_{p})^{-1}{\bm{\Sigma}}\right],$
for $z\in\mathbb{C}/\mathbb{R}^{+}$. Using averaged local law Knowles & Yin (2017), we obtain
$\frac{1}{n}\operatorname{Tr}\left[(n^{-1}\mathbf{X}^{\top}\mathbf{X}-zI_{p})^{-1}{\bm{\Sigma}}\right]$ $\overset{\textrm{a.s.}}{\longrightarrow}\frac{-1}{z}\,\gamma\int\frac{x}{mx+1}dH(x),$
@@ A.3 Bias
$\mathbb{E}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]=(\widetilde{w}A_{\lambda}M)^{t}A_{\lambda}M\boldsymbol{\beta}+w\sum_{i=0}^{t-1}(\widetilde{w}A_{\lambda}M)^{i}A_{\lambda}M\boldsymbol{\beta}$ (A.6)
@@
$=A_{\lambda}M\boldsymbol{\beta}$
@@
$\mathbb{E}[\widehat{\boldsymbol{\beta}}_{k+1,\lambda}|\mathbf{X}]$ $=A_{\lambda}M(w\boldsymbol{\beta}+(1-w)\mathbb{E}[\widehat{\boldsymbol{\beta}}_{k,\lambda}|\mathbf{X}])$
$=wA_{\lambda}M\boldsymbol{\beta}+\widetilde{w}A_{\lambda}M\mathbb{E}[\widehat{\boldsymbol{\beta}}_{k,\lambda}|\mathbf{X}])$
$=wA_{\lambda}M\boldsymbol{\beta}+\widetilde{w}A_{\lambda}M\left[(\widetilde{w}A_{\lambda}M)^{k}A_{\lambda}M\boldsymbol{\beta}+w\sum_{i=0}^{k-1}(\widetilde{w}A_{\lambda}M)^{i}A_{\lambda}M\boldsymbol{\beta}\right]$
$=(\widetilde{w}A_{\lambda}M)^{k+1}A_{\lambda}M\boldsymbol{\beta}+w\sum_{i=0}^{k}(\widetilde{w}A_{\lambda}M)^{i}A_{\lambda}M\boldsymbol{\beta}$
@@
$B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\left(\frac{n\lambda}{w}\right)^{2}\|A_{\lambda/w}(I-(\widetilde{w}A_{\lambda}M)^{t+1})\boldsymbol{\beta}\|_{\bm{\Sigma}}^{2}$ (A.7)
$\lim_{t\to\infty}B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\left(\frac{n\lambda}{w}\right)^{2}\|A_{\lambda/w}\boldsymbol{\beta}\|_{\bm{\Sigma}}^{2}=\left(\frac{n\lambda}{w}\right)^{2}\operatorname{Tr}[\boldsymbol{\beta}\boldsymbol{\beta}^{\top}A_{\lambda/w}{\bm{\Sigma}}A_{\lambda/w}]$ (A.8)
@@
$\boldsymbol{\beta}-\mathbb{E}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]=\,\,$ $\boldsymbol{\beta}-(\widetilde{w}A_{\lambda}M)^{t}A_{\lambda}M\boldsymbol{\beta}$
$-w(I-\widetilde{w}A_{\lambda}M)^{-1}(I-(\widetilde{w}A_{\lambda}M)^{t})A_{\lambda}M\boldsymbol{\beta}$ (A.9)
Diagonalizing $A_{\lambda}M$ and and simplifying using the following identity
@@
$\boldsymbol{\beta}-\mathbb{E}[\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X}]=(I-A_{\lambda}M)(I-\widetilde{w}A_{\lambda}M)^{-1}(I-(\widetilde{w}A_{\lambda}M)^{1+t})\boldsymbol{\beta}$
@@
to conclude $(I-A_{\lambda}M)(I-\widetilde{w}A_{\lambda}M)^{-1}=\frac{n\lambda}{w}A_{\lambda/w}$. This completes the proof of equation (A.7). equation (A.8) follows since $\sigma_{\max}(\widetilde{w}A_{\lambda}M)<\widetilde{w}<1$. ∎
@@
$z^{2}\operatorname{Tr}[CQ_{n}(z){\bm{\Sigma}}Q_{n}(z)]\sim\operatorname{Tr}[C(I+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\frac{1}{1-\frac{1}{n}df_{2}(1/m(z))}$ (A.10)
where $df_{2}(\kappa)=\operatorname{Tr}[{\bm{\Sigma}}^{2}(\kappa I+{\bm{\Sigma}})^{-2}]$ and $Q_{n}(z)=(n^{-1}M-zI)^{-1}$.

### v2

@@ 4 Simulations
Figure 2: In $(a)$ and $(b)$, we plot the optimal mixing weight $w^{\star}$ as a function of $\lambda$ for different values of $\gamma$ and two classes of covariance matrices: Panel (a) considers ${\bm{\Sigma}}=\bm{I}$ with high noise $\sigma^{2}=64$, demonstrating $w^{\star}$ can be close to $0.5$ for low SNR. Panel $(b)$ plots $w^{\star}$ for the spiked covariance matrix showing $w^{\star}=1$ for large $\lambda$. Panel $(c)$ plots the generalization error as a function of $\lambda$ for $\bm{\Sigma}$} equicorrelated matrix. Here, empirical risks align with theoretical predictions given by Proposition 3.1, though $\bm{\Sigma}$} violates Assumption 2.1, illustrating the robustness of our results.
Figure 3: In $(a)$, we plot the optimal $w^{\star}$ for ridge regression as a function of $\gamma$. In $(b)$, we plot the empirical risk for the min-norm interpolator to demonstrate the dynamic mixing scenario described in Section 3.1. Note that, the risk under dynamic mixing and $w=\varphi^{-1}$ heavily overlap for $t=2$ onward. In $(c)$, we plot the empirical (the points) & theoretical (the solid lines) risks for the no fresh data augmentation case described by equation (5.1) and Theorem 5.1. Note the risk is optimized at $w^{\star}=1$. Throughout ${\bm{\Sigma}}=\bm{I}$.
5 Beyond fresh data augmentation
In this section, we provide rigorous evidence on when and how model collapse can be prevented beyond the fresh data augmentation framework equation (1). In the interest of space, we focus on the min-$\ell_{2}$-norm interpolator, but our results can be extended for ridge regression.
No fresh real data First, we consider the setup where fresh real data is unavailable beyond the original data $(\mathbf{y},\mathbf{X})$. Therefore, one augments the initial labels $\mathbf{y}$ with synthetic labels as follows
$\widehat{\boldsymbol{\beta}}^{(r)}_{t}:=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}\big(w\mathbf{y}+(1-w)\widetilde{\mathbf{y}}_{t}\big)\,\,\text{where}\,\,\widetilde{\mathbf{y}}_{t}=\mathbf{X}\widehat{\boldsymbol{\beta}}^{(r)}_{t-1}+\widetilde{\bm{\varepsilon}}_{t}.$ (5.1)
We assume $\widehat{\boldsymbol{\beta}}^{(r)}_{0}$ is the min-$\ell_{2}$-norm interpolator calculated using the original $(\mathbf{X},\mathbf{y})$.
Theorem 5.1.
The risk of $\widehat{\boldsymbol{\beta}}^{(r)}_{t}$, defined by  equation (5.1), satisfies the following. For any $w\in(0,1)$, we have almost surely over the randomness in the covariates,
$\lim_{n\to\infty}\lim_{t\to\infty}R(\widehat{\boldsymbol{\beta}}^{(r)}_{t};\boldsymbol{\beta})=\sigma^{2}\frac{1}{w(2-w)}\mathcal{V}+b_{\star}\mathcal{B}$ (5.2)
with $\mathcal{V},\mathcal{B}$ defined as in equation (3.5). The limiting risk is minimized at $w^{\star}=1$.
In Theorem 5.1, $\mathcal{V},\mathcal{B}$ do not depend on the weight, thus the asymptotic risk is monotonically decreasing in $w$, yielding $w^{\star}=1$. Thus, the asymptotic risk is minimized when the weight on the real data is $1$, suggesting adding synthetic data following this mechanism fails to improve performance. This can be seen empirically in Figure 3 (c), where we plot the empirical risk as well as the theoretical risk. Here $n=500$, $\texttt{SNR}=1$ and we consider $50$ iterations.
Time-varying covariates Next, we consider the setup where different covariates are used to train different iterations of the model. Formally, for $t\geq 1$, we have real labels: $\mathbf{y}_{t}\leftarrow\mathbf{X}_{t}\boldsymbol{\beta}+{\bm{\varepsilon}}_{t}$ as well as synthetic labels: $\widetilde{\mathbf{y}}_{t}\leftarrow\mathbf{X}_{t}\widehat{\boldsymbol{\beta}}_{t-1}+\widetilde{\bm{\varepsilon}}_{t}$. The estimator for $t$-th iteration is defined as
$\widehat{\boldsymbol{\beta}}^{(v)}_{t}:=(\mathbf{X}^{\top}_{t}\mathbf{X}_{t})^{\dagger}\mathbf{X}^{\top}_{t}\big(w\mathbf{y}_{t}+(1-w)\widetilde{\mathbf{y}}_{t}\big).$ (5.3)
We provide the precise generalization error of $\widehat{\boldsymbol{\beta}}^{(v)}_{t}$ in Theorem E.1 (for simplicity we assume ${\bm{\Sigma}}=\bm{I}$, see also Remark E.1 regarding limit orders) and characterize the unique optimal weight at which the risk is minimized. Unlike Theorem 3.1, both the bias and variance depend on $w$.
Time-varying covariates with fresh labels in a fraction Finally, we consider a setting where at each iteration, fresh labels are available only for a fraction of covariates. Since all covariates do not both see real and synthetic labels, instead of our mixing framework from equation (1), we consider pooling the datasets with real and synthetic labels. Formally, for $t\geq 1$, we observe real responses: $\mathbf{y}_{t}\leftarrow\mathbf{X}_{t}\boldsymbol{\beta}+{\bm{\varepsilon}}_{t}$ and synthetic responses: $\widetilde{\mathbf{y}}_{t}\leftarrow\widetilde{\mathbf{X}}_{t}\widehat{\boldsymbol{\beta}}_{t-1}+\widetilde{\bm{\varepsilon}}_{t}$. Define
$\widehat{\boldsymbol{\beta}}^{(a)}_{t}\leftarrow(\mathbf{X}^{\top}_{t}\mathbf{X}_{t}+\widetilde{\mathbf{X}}^{\top}_{t}\widetilde{\mathbf{X}}_{t})^{\dagger}\big(\mathbf{X}^{\top}_{t}\mathbf{y}_{t}+\widetilde{\mathbf{X}}^{\top}_{t}\widetilde{\mathbf{y}}_{t}\big).$
In Appendix E.3 (specifically Theorem E.2), we show that there exist ranges of values of $\gamma$ where preventing model collapse is not possible in this setting.
@@ 6 Discussion
We provide a rigorous analysis of model collapse under overparametrization for linear models. As an overarching theme, we demonstrate how mixing real-data with synthetic outputs mitigates model collapse, and identify optimal mixing ratios that minimize prediction error in this context. As a promising next direction, understanding how model collapse affects interpolators in other $\ell_{p}$ geometries would be crucial. Such interpolators arise as implicit regularized limits of popular algorithms Gunasekar et al. (2018a) and typically require techniques beyond random matrix theory — a critical tool employed in our analysis. Additionally, approaches that extend linear arguments to non-linear high-dimensional problems (c.f., Sur and Candès (2019); Hu and Lu (2022)) should enable our qualitative conclusions to generalize to structured non-linear models. This includes generalized linear models, single-index models, and even non-parametric models through parametric-to-non-parametric equivalence techniques introduced in Lahiry and Sur (2023). Finally, extending our results to more complex architectures remains an important challenge. One promising approach involves studying multi-index models and their sequential variants, which capture classes of neural networks and transformer architectures Troiani et al. (2024; 2025); Cui (2025). Investigating model collapse and mitigation strategies in these contexts presents an exciting avenue for future research.
Supplementary Material
@@ A.1 Notations
$\mathbf{M}:=\mathbf{X}^{\top}\mathbf{X},\quad\quad\mathbf{A}_{\lambda}:=(\mathbf{M}+n\lambda\bm{I})^{-1}.$
Note that $\mathbf{A}_{\lambda}$ and $\mathbf{M}$ are simultaneously diagonalizable and thus commute with each other. If $\lambda=0$, we use the notation $\mathbf{A}_{0}=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$. The mixing proportion of the real-data is denoted by $w$ and we define $\widetilde{w}:=1-w$. For two sequence of random variables $u_{n},v_{n}$, we use the notation $u_{n}\sim v_{n}$ if $u_{n}/v_{n}\xrightarrow{P}1$ as $n\rightarrow\infty$.
@@ A.2 Variance
$\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2(k-1)}\mathbf{M}^{2k-1}\mathbf{A}_{\lambda}^{2k}+\widetilde{w}^{2t}\mathbf{M}^{2t+1}\mathbf{A}_{\lambda}^{2t+2}\right]$ (A.1)
@@
We will prove equation (A.1) by induction. For the base case $(t=0)$, we have $\widehat{\boldsymbol{\beta}}_{0,\lambda}=\mathbf{A}_{\lambda}\mathbf{X}^{\top}\mathbf{y}$. Hence,
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{0,\lambda}|\mathbf{X}]$ $=\mathbf{A}_{\lambda}\mathbf{X}^{\top}{\sf Cov}[\mathbf{y}|\mathbf{X}]\mathbf{X}\mathbf{A}_{\lambda}$
$=\mathbf{A}_{\lambda}\mathbf{X}^{\top}(\sigma^{2}\bm{I})\mathbf{X}\mathbf{A}_{\lambda}=\sigma^{2}\mathbf{A}_{\lambda}^{2}\mathbf{M}$
@@
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t+1,\lambda}|\mathbf{X}]$ $=\mathbf{A}_{\lambda}\mathbf{X}^{\top}{\sf Cov}[w\mathbf{y}_{t}+\widetilde{w}\widetilde{\mathbf{y}}_{t}|\mathbf{X}]\mathbf{X}\mathbf{A}_{\lambda}$
$=\mathbf{A}_{\lambda}\mathbf{X}^{\top}(w^{2}\sigma^{2}I+\widetilde{w}^{2}\mathbf{X}{\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]\mathbf{X}^{\top}+\sigma^{2}\widetilde{w}^{2}I)\mathbf{X}\mathbf{A}_{\lambda}$
$=\widetilde{w}^{2}\mathbf{A}_{\lambda}\mathbf{M}{\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]\mathbf{M}\mathbf{A}_{\lambda}+\sigma^{2}(w^{2}+\widetilde{w}^{2})\mathbf{M}\mathbf{A}_{\lambda}^{2}$
$=\widetilde{w}^{2}\mathbf{A}_{\lambda}\mathbf{M}\left(\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2(k-1)}\mathbf{M}^{2k-1}\mathbf{A}_{\lambda}^{2k}+\widetilde{w}^{2t}\mathbf{M}^{2t+1}\mathbf{A}_{\lambda}^{2t+2}\right]\right)\mathbf{M}\mathbf{A}_{\lambda}$
$\quad\quad+\sigma^{2}(w^{2}+\widetilde{w}^{2})\mathbf{M}\mathbf{A}_{\lambda}^{2}$
$=\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2k}\mathbf{M}^{2k+1}\mathbf{A}_{\lambda}^{2k+2}+\widetilde{w}^{2t+2}\mathbf{M}^{2t+3}\mathbf{A}_{\lambda}^{2t+4}\right]$
$\quad\quad+\sigma^{2}(w^{2}+\widetilde{w}^{2})\mathbf{M}\mathbf{A}_{\lambda}^{2}$
$=\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t+1}\widetilde{w}^{2(k-1)}\mathbf{M}^{2k-1}\mathbf{A}_{\lambda}^{2k}+\widetilde{w}^{2t+2}\mathbf{M}^{2t+3}\mathbf{A}_{\lambda}^{2t+4}\right]$
@@
$V(\boldsymbol{\beta}_{t,\lambda};\boldsymbol{\beta})$ $=\sigma^{2}\left[(w^{2}+\widetilde{w}^{2})\sum_{k=1}^{t}\widetilde{w}^{2(k-1)}\operatorname{Tr}[\mathbf{B}_{\lambda}^{k}]+\widetilde{w}^{2t}\operatorname{Tr}[\mathbf{B}_{\lambda}^{t+1}]\right]$ (A.2)
where $\mathbf{B}_{\lambda}^{k}=\mathbf{M}^{2k-1}\mathbf{A}_{\lambda}^{2k}{\bm{\Sigma}}$.
Next, note that $\sigma_{\max}(\mathbf{M}\mathbf{A}_{\lambda})<1$ and therefore ${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]$ converges in matrix norm to the following limit:
$\lim_{t\to\infty}{\sf Cov}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]$ $=\sigma^{2}(w^{2}+(1-w)^{2})\sum_{k=0}^{\infty}(1-w)^{2k}(\mathbf{X}^{\top}\mathbf{X})^{2k+1}\mathbf{A}_{\lambda}^{2k+2}$
$=\sigma^{2}(w^{2}+(1-w)^{2})\mathbf{X}^{\top}\mathbf{X}\mathbf{A}_{\lambda}^{2}(I-(1-w)^{2}(\mathbf{X}^{\top}\mathbf{X})^{2}A_{\lambda}^{2})^{-1}$
@@
$=\frac{\sigma^{2}(w^{2}+(1-w)^{2})}{2(1-w)}\left[(w\mathbf{X}^{\top}\mathbf{X}+n\lambda\bm{I})^{-1}-((2-w)\mathbf{X}^{\top}\mathbf{X}+n\lambda\bm{I})^{-1}\right]$
$\implies\lim_{t\to\infty}V(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\frac{\sigma^{2}(w^{2}+(1-w)^{2})}{2(1-w)}\left(\frac{1}{w}\operatorname{Tr}[\mathbf{A}_{\lambda/w}{\bm{\Sigma}}]-\frac{1}{2-w}\operatorname{Tr}[\mathbf{A}_{\lambda/(2-w)}{\bm{\Sigma}}]\right)$
@@
$\lim_{n\to\infty}\operatorname{Tr}[\mathbf{A}_{-z}{\bm{\Sigma}}]=\lim_{n\to\infty}\frac{1}{n}\operatorname{Tr}\left[(n^{-1}\mathbf{X}^{\top}\mathbf{X}-z\bm{I}_{p})^{-1}{\bm{\Sigma}}\right],$
for $z\in\mathbb{C}/\mathbb{R}^{+}$. Using averaged local law Knowles and Yin (2017), we obtain
$\frac{1}{n}\operatorname{Tr}\left[(n^{-1}\mathbf{X}^{\top}\mathbf{X}-z\bm{I}_{p})^{-1}{\bm{\Sigma}}\right]$ $\overset{\textrm{a.s.}}{\longrightarrow}\frac{-1}{z}\,\gamma\int\frac{x}{mx+1}dH(x),$
@@ A.3 Bias
$\mathbb{E}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]=(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{t}\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}+w\sum_{i=0}^{t-1}(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{i}\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}$ (A.6)
@@
$=\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}$
@@
$\mathbb{E}[\widehat{\boldsymbol{\beta}}_{k+1,\lambda}|\mathbf{X}]$ $=\mathbf{A}_{\lambda}\mathbf{M}(w\boldsymbol{\beta}+(1-w)\mathbb{E}[\widehat{\boldsymbol{\beta}}_{k,\lambda}|\mathbf{X}])$
$=w\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}+\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M}\mathbb{E}[\widehat{\boldsymbol{\beta}}_{k,\lambda}|\mathbf{X}])$
$=w\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}+\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M}\left[(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{k}\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}+w\sum_{i=0}^{k-1}(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{i}\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}\right]$
$=(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{k+1}\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}+w\sum_{i=0}^{k}(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{i}\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}.$
@@
$B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\left(\frac{n\lambda}{w}\right)^{2}\|\mathbf{A}_{\lambda/w}(\bm{I}-(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{t+1})\boldsymbol{\beta}\|_{\bm{\Sigma}}^{2}.$ (A.7)
$\lim_{t\to\infty}B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\left(\frac{n\lambda}{w}\right)^{2}\|\mathbf{A}_{\lambda/w}\boldsymbol{\beta}\|_{\bm{\Sigma}}^{2}=\left(\frac{n\lambda}{w}\right)^{2}\operatorname{Tr}[\boldsymbol{\beta}\boldsymbol{\beta}^{\top}\mathbf{A}_{\lambda/w}{\bm{\Sigma}}\mathbf{A}_{\lambda/w}].$ (A.8)
@@
$\boldsymbol{\beta}-\mathbb{E}[\widehat{\boldsymbol{\beta}}_{t,\lambda}|\mathbf{X}]=\,\,$ $\boldsymbol{\beta}-(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{t}\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}$
$-w(I-\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{-1}(I-(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{t})\mathbf{A}_{\lambda}\mathbf{M}\boldsymbol{\beta}.$ (A.9)
Diagonalizing $\mathbf{A}_{\lambda}\mathbf{M}$ and simplifying using the following identity
@@
$\boldsymbol{\beta}-\mathbb{E}[\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X}]=(\bm{I}-\mathbf{A}_{\lambda}\mathbf{M})(\bm{I}-\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{-1}(\bm{I}-(\widetilde{w}A_{\lambda}\mathbf{M})^{1+t})\boldsymbol{\beta}$
@@
to conclude $(\bm{I}-\mathbf{A}_{\lambda}\mathbf{M})(\bm{I}-\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})^{-1}=\frac{n\lambda}{w}\mathbf{A}_{\lambda/w}$. This completes the proof of equation (A.7). equation (A.8) follows since $\sigma_{\max}(\widetilde{w}\mathbf{A}_{\lambda}\mathbf{M})<\widetilde{w}<1$. ∎
@@
$z^{2}\operatorname{Tr}[\bm{C}\bm{Q}_{n}(z){\bm{\Sigma}}\bm{Q}_{n}(z)]\sim\operatorname{Tr}[\bm{C}(\bm{I}+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\frac{1}{1-\frac{1}{n}df_{2}(1/m(z))}$ (A.10)
where $df_{2}(\kappa)=\operatorname{Tr}[{\bm{\Sigma}}^{2}(\kappa\bm{I}+{\bm{\Sigma}})^{-2}]$ and $\bm{Q}_{n}(z)=(n^{-1}\mathbf{M}-z\bm{I})^{-1}$.